OpenStudy (anonymous):

GIVING MEDAL !!!!!!!!! Choose the single logarithmic expression to 2 log 4 + log 2 +log 2

OpenStudy (anonymous):

2 log 4 + log 2 +log 2 = \[ 2\log (2)^2 + \log 2 +\log 2= 2\times 2\log 2 + \log 2 +\log 2= 4\log 2 + \log 2 +\log 2=6\log2\]